OpenStudy (anonymous):

Given the quadratic function f(x) = x2 – 12x + 36, find a value of x such that f(x) = 25

OpenStudy (chaise):

The answer is 1. x^2 -12x+36 = 25 x^2-12x= -11 x(x-12) = -11 x = 0 OR x-12=-11 x = 1 Disregard the 0 answer.

OpenStudy (anonymous):

Shouldn't the answers be 1 union 11?

OpenStudy (anonymous):

x^2 -12x+36 = 25 x^2-12x+11=0 (x-11)(x-1)=0 x=11 or x=1